OpenStudy (anonymous):

Geometry The Area Of A Rectangle Is At Most 21 Square Inches. The Width Of The Rectangle Is 3.5. What Are The Possible Measurements For The Length Of The Rectangle.

OpenStudy (anonymous):

A = l * w 21 = L * 3.5 simplify